Shift report 20011007 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Oct 07 2001 - 00:08:14 EDT

  • Next message: Apache: "Shift report 20011007 00:00-08:00"

    ------------------------------------------------------------
    Date: Sun Oct  7 00:08:14 2001, Shift: 16:00-24:00
    Supervisor: hiro, with: 
    Log book#9 pages 107 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    1.  MC is studying beta*=2m for an entire shift
       Beam is stored, but not steered for our IR.
    
    2.  NO runs.
    3.  None detected.
       During the shift, around 8:30PM, the wirewall equipment
    in the little room at counting house beepped for about 2 
    minutes.  I do not know if this is normal.  
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Oct 07 2001 - 00:08:55 EDT